5. (-/1 Points) DETAILS TANAPCALC10 6.5.008.MI. MY NOTES Evaluate the definite integral. dx Need Help? Read Talk to a Tutor D please circle answers thank you!
Answer #1

Solution ndn n +7 Substitute nt ndnat at x²+7=t² 4 t at -) 1 t 7 bort - I - J duc r7 -) I tu ſ7 bet 901 -7 | I = 4-57 G